# Direction Symmetry of Wave Field Modulation by Tidal Current

### v1

@@ Abstract
Theoretical studies on the modulation of unidimensional regular waves over a flat bottom due to a current typically assign an asymmetry between the effects of opposing/following streams on the evolution of major sea variables, such as significant wave height. The significant wave height is expected to monotonically increase with opposing streams and to decrease with following streams. To some extent, observations on data sets containing a few thousand of waves or over a continuous series of about a day confirm this prediction. Here, based on a multi-year dataset, we show that in very broad-banded seas with high directional spread, the asymptotic behavior of sea variables is highly non-trivial and does not follow the theoretical predictions, especially at high values of the ratio between tidal stream and group speed.
@@ I INTRODUCTION
The deterministic study of inhomogeneous waves sprung near the middle of the twentieth century [1, 2]. Advanced mathematical techniques arose in the following decades [3, 4, 5, 6], in particular with wave-current interaction being assessed through ray theory [7, 8], linear wave theory [9, 10], radiation stress [11, 12], spectral [13] and perturbative methods [14]. Through these theoretical advances, transformations of wave characteristics were measured and confirmed in major global currents, such as the Agulhas current [15], the Gulf Stream [16, 17] and the Kuroshio [18, 19]. These transformations concern both significant wave height and wave energy, as well as spectral bandwidth and directional spreading. Additionally, the transformations have been corroborated in laboratory experiments [20, 21, 22, 23, 24]. The first studies on wave-current systems focused on tides [1, 25, 26], theoretically treating them as steady currents. They found that currents either reduce or expand the wavelength, whose change is met with a modulation in the wave height [1]. These properties of the wave-current system seem to be universal, featuring similar changes in wave height, length, and period in estuaries and rivers [27, 28, 29]. However, simplified quasi-stationary or quasi-homogeneous models based on the energy balance equations underestimate the effect of tidal currents on wave heights [30, 31, 32].
@@
The majority of the mentioned studies limit the wave-current system to waves encountering an opposing current, resulting in a shortening and steepening of waves due to energy focusing. Only a few studies considered the case of waves propagating in the same direction as the current, and their findings were ambiguous. Initial theoretical considerations suggested that the wavelength should increase when waves propagate in the same direction as a tidal stream [1, 26]. Indeed, MacIver et al. [23] reported from an experimental study that in following currents waves become longer, while at the same time their height is reduced. They concluded that in waves with a following current, wave energy is decreased due to the strain rate of the current speed. Swan et al. [22], however, found in an experimental study, supported by a numerical model, that the crest-trough asymmetry of waves propagating in a following current increases towards higher and sharper wave crests. This asymmetry should increase the probability of extreme waves [33, 34, 35, 36, 37, 38, 39].
Observational studies to date are mostly focused on deterministic investigations for specific events or short-term time series analysis [40, 41, 42, 43, 44], even those with larger data sets [45, 46, 47, 48]. Here, we investigate the behavior of the modulation due to the wave-current interaction of long-term series from a statistical point of view, namely of average wave properties. We characterize the symmetry in the evolution of fundamental wave properties in response to opposing and following currents. We show that following currents can amplify wave fields in deep water in addition to coastal waters [43] and that both the nonlinearity of the sea state and the speed of the tidal stream impact the magnitude and asymmetry of the modulation.
@@ II Data and methods
Wave elevation data were recorded by a surface-following directional Waverider buoy of type MkIII. The sampling frequency was 1.28 Hz. The resolution of the wave buoy is specified as 0.01 m, with an accuracy of less than 0.5% of the measured surface elevation relative to the calibrated still water line [51]. The wave buoy delivered surface elevation data in samples of 30 minutes in length, containing 366 waves on average. The samples were quality-controlled according to the procedure described in Teutsch et al. [52] and subsequently used to calculate the directional spectra (figure 2a,b) and the sea state parameters in table 1.
@@
Group speed $c_{\mathrm{g}}=\frac{1}{2}\sqrt{\frac{g}{k_{\mathrm{p}}}\tanh(k_{\mathrm{p}}h)}\cdot(1+\frac{2k_{\mathrm{p}}h}{\sinh(2k_{\mathrm{p}}h)})$
@@
Steepness $\varepsilon=\frac{\sqrt{2}}{\pi}k_{p}H_{s}$
Table 1: Sea parameter definitions, with gravity $g$ and spectral moments $m_{n}$.
@@
Figure 4: Histograms of the frequency of 30-minute wave records as a function of the major dimensional variables as well as normalized current speed.
@@
with $C_{1}=\sqrt{a_{1}^{2}+b_{1}^{2}}$, in which $a_{1}$ and $b_{1}$ denote the first sine and cosine Fourier coefficients of the directional spectrum, respectively. The significant wave height $H_{s}$ was computed as the mean of the highest third of waves in each 30-minute sample. To evaluate the effect of current speed $U$ and current direction on major sea state parameters, we selected two sets of samples, for which the angle between peak wave direction and current direction was either $0\pm 10^{\circ}$ (henceforth referred to as ”following current”) or $180\pm 10^{\circ}$ (henceforth referred to as ”opposing current”). Within these ranges, sine values keep below 0.17 and cosine values stay beyond ±$\pm 0.98$, so that the transverse flow is negligible and the axial current is virtually unaffected. References without current, hereafter denoted rest conditions, were defined as samples with $|U/c_{g}|<0.02$, $c_{g}$ being the group speed of the peak frequency of the wave spectrum. This threshold is consistent with the residual transverse current at current incidence angles of ±$\pm 10^{\circ}$.
@@
In total, the quality-controlled data set contained 4686 samples (2156 with opposing current, 2321 with forward current, 209 in rest conditions). Values of $U/c_{g}$ were binned in intervals of $\sim 1/50$, except around zero, where a single bin covers $|U/c_{g}|\leq 0.02$, and for the outermost bins, where data are sparser.
@@
The histograms in figure 4 provide descriptive statistics of all selected 30 minute samples along the entire data set. Figures 4a,d describe the near-symmetrical character of the tidal speeds, whether in dimensional units or normalized by the peak group speed. Figures 4b,c provide the wave climate of the region, with mild sea characteristics for both significant wave height and zero-crossing period. Most of the data is in intermediate ($k_{p}h=0.3-3.1$, 47% of the dataset) to deep water ($k_{p}h>\pi$, 50% of the dataset, or even 79% corresponding to $k_{p}h\geq 2$) (figure 4e). Adding to the complexity of the sea conditions, the spectrum is significantly directionally spread (figure 4f). Furthermore, the spectrum is very broad (Figure 4h). The sea is composed of waves typically of second order in wave steepness ($0.05\leq\varepsilon\leq 0.09$) [57], albeit smaller fractions of the dataset are also of first and higher orders (figure 4g). As discussed in the next section, this strong nonlinearity will be key to the physical behavior of the waves. Typically, consideration of the nonlinear regime of the wave steepness and its effect on wave-tide modulation is overlooked in most studies.
III Results
Figure 5 displays the dependence of the main wave parameters on the relative current velocity. Negative current velocity refers to a current opposing the peak wave direction. Only the mean wave steepness (figure 5a) displays a clear asymmetry between forward and opposing currents. In spite of the wide absolute confidence ranges, the size of the samples is quite large, typically with 100,000 entries per bin. As such, pairwise Welch’s $t$-tests [58] show a highly significant difference ($p<10^{-3}$) between bins corresponding to forward and opposing currents of the same speed, ±$\pm U/c_{g}$, for all pairs of bins.
Other variables display a much more symmetric behavior. The anticipated increase in significant wave height due to an opposing current [1, 10] is observed in our data (figure 5b), but restricted to $|U/c_{g}|\leq 0.05$. Beyond, this trend is inverted and $H_{s}$ decreases again. This decrease might be due to either wave blocking or increasing dissipation due to high mean steepness, although it seems to appear earlier than expected by Toffoli et al. [59]. The behavior is comparable to the effect of forward currents: $H_{s}$ increases for mild current speeds, unlike the expected behavior of regular waves [23]. As in the case of an opposing stream, the significant wave height decreases for $U/c_{g}>0.05$, likely due to wave dissipation.
@@ III Results
From the classical literature [10], we know that the dispersion relation of narrow-banded, unidirectional, and linear waves is affected by the presence of a current: The dispersion relation $\omega_{0}^{2}=gk_{0}$ becomes $(\omega\pm Uk)^{2}=gk_{0}$. Bearing in mind that in deep water $\omega/k=c_{p}=2c_{g}$, with $c_{p}$ denoting the phase velocity, one finds the modulation of the wave number to be of the order of $k/k_{0}\approx 1/[1\pm(U/2c_{g})]^{2}\approx 1\mp(U/c_{g})+3/4(U/c_{g})^{2}$. When $|U/c_{g}|\ll 0.1$, this modulation is expected to be strongly asymmetrical, however larger values of $U/c_{g}$ will diminish such asymmetry, as the quadratic term increases and tends to dominate. Furthermore, the wave theory for steeper waves leads to a strengthening of the quadratic term [60, 61, 62], in addition to shear terms [63, 22]. Besides, from the dynamic point of view, Ho et al. [43] showed that the relevant parameter is the phase speed of the current relative to the wave group speed, so that the interaction between a following current and a comparatively slow wave field in the fixed frame appears to be an interaction between an opposing current and waves in the reference frame of the tide.

### v2

@@ Abstract
Theoretical studies on the modulation of unidimensional regular waves over a flat bottom due to a current typically assign an asymmetry between the effects of opposing/following streams on the evolution of major sea variables, such as significant wave height. The significant wave height is expected to monotonically increase with opposing streams and to decrease with following streams. To some extent, observations on data sets containing a few thousand of waves or over a continuous series of about a day confirm this prediction. Based on a multi-year dataset, we show that in very broad-banded seas with high directional spread especially at high values of the ratio between tidal stream and group velocity, the asymptotic behavior of sea variables is highly non-trivial and does not follow the theoretical predictions, thus featuring a symmetrical modulation of several sea state variables respective to whether the current opposes or follows the direction of wave motion.
@@ I INTRODUCTION
The deterministic study of inhomogeneous waves sprung near the middle of the twentieth century [1, 2]. Advanced mathematical techniques arose in the following decades [3, 4, 5, 6], in particular with wave-current interaction being assessed through ray theory [7, 8], linear wave theory [9, 10], radiation stress [11, 12], action balance [13], spectral [14] and perturbative methods [15]. Through these theoretical advances, transformations of wave characteristics were measured and confirmed in major global currents, such as the Agulhas current [16], the Gulf Stream [17, 18] and the Kuroshio [19, 20]. These transformations concern both significant wave height and wave energy, spectral bandwidth and directional spreading. Additionally, the transformations have been corroborated in laboratory experiments [21, 22, 23, 24, 25]. The first studies on wave-current systems focused on tides [1, 26], finding that currents either reduce or expand the wavelength, whose change is met with a modulation in the wave height [1]. These properties of the quasi-stationary and/or homogeneous wave-current systems seem to be universal, featuring similar changes in wave height, length, and period in estuaries and rivers [27, 28, 29]. On the other hand, Barber [30] was one of the first to address currents as dominantly changing in time and imply that the absolute period of a near-monochromatic swell is modified following the Doppler effect due to currents, but that the wavelength remains unchanged. As pointed out by Tolman [31], simplified quasi-stationary or quasi-homogeneous theoretical models based on the action balance equations underestimate the effect of tidal currents on wave heights. Furthermore, Tolman [32] and Holthuijsen and Tolman [17] studied the problem with a full non-stationary and non-homogeneous approach, showing non-local aspects of wave-current interactions in the Southern North Sea and Gulf Stream, respectively.
@@
The majority of the mentioned studies limit the wave-current system to waves encountering an opposing current, resulting in a shortening and steepening of waves due to energy focusing. Only a few studies considered the case of waves propagating in the same direction as the current, and their findings were ambiguous. Theoretical considerations suggested that the wavelength should increase when waves propagate in the same direction as a tidal stream [1]. Indeed, MacIver et al. [24] reported from an experimental study that in following currents waves become longer, while at the same time their height is reduced. They concluded that in waves with a following current, wave energy is decreased due to the strain rate of the current speed. Swan et al. [23], however, found in an experimental study, supported by a numerical model, that the crest-trough unevenness of waves propagating in a following current increases towards higher and sharper wave crests. This unevenness should increase the probability of extreme waves [33, 34, 35, 36, 37, 38, 39].
Observational studies to date are mostly focused on deterministic investigations for specific events or short-term time series analysis [40, 41, 42, 43, 44], even those with larger data sets [45, 46, 47, 48]. Here, we investigate the behavior of the modulation due to the wave-current interaction of long-term series from a statistical point of view, namely of average wave properties. We characterize the symmetry in the evolution of fundamental wave properties in response to opposing and following currents in the limited context of currents being co-linear to the direction of wave motion. We show that the following currents can amplify wave fields in deep water in addition to coastal waters [43] and that both the nonlinearity of the sea state and the speed of the tidal stream impact the magnitude and asymmetry of the modulation.
@@ II Data and methods
group velocity $c_{\mathrm{g}}=\frac{1}{2}\sqrt{\frac{g}{k_{\mathrm{p}}}\tanh(k_{\mathrm{p}}h)}\cdot(1+\frac{2k_{\mathrm{p}}h}{\sinh(2k_{\mathrm{p}}h)})$
@@
Steepness $\varepsilon=k_{p}H_{s}$
Table 1: Sea parameter definitions, with gravity $g$ and spectral moments $m_{n}$ assuming small current effects.
Figure 4: Histograms of the frequency of 30-minute wave records as a function of the major dimensional variables, as well as normalized current speed. Brighter colors emphasize higher probability density.
Wave elevation data were recorded by a surface-following directional Waverider buoy of type MkIII. The sampling frequency was 1.28 Hz. The resolution of the wave buoy is specified as 0.01 m, with an accuracy of less than 0.5% of the measured surface elevation relative to the calibrated still water line [51]. The wave buoy delivered surface elevation data in samples of 30 minutes in length, containing 366 waves on average. The samples were quality-controlled according to the procedure described in Teutsch et al. [52] and subsequently used to calculate the directional spectra (figure 2a,b) and the sea state parameters in table 1 for the finite water depth water waves without current effects. The current effect on dispersion has ramifications to all other variables, and reads (bold denotes vectors) [11]:
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\omega=\omega_{0}+\textbf{U}\cdot\textbf{k}\quad,\quad\omega_{0}=g|\textbf{k}|\tanh{|\textbf{k}|h}}$ (1)
thus leading to,
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\omega^{2}_{0}=\omega^{2}\left(1-\frac{\textbf{U}\cdot\textbf{k}}{\omega}\right)^{2}=g|\textbf{k}|\tanh{|\textbf{k}|h}}$ (2)
As observed in figure 4, most of the data lies in the deep water regime, such that the hyperbolic function has small impact on the dispersion. Moreover, the typical values of the variables measured by their peak parameters show that $U\sim\pm 0.5$m/s, has periods of $T_{p}\sim 6$s and length $\lambda_{p}\sim 75$m and therefore the correction due to the current on the dispersion is small $\textbf{U}\cdot\textbf{k}/\omega\sim UT_{p}/\lambda_{p}\approx 0.04\ll 1$. Therefore, we can compute peak spectral parameters without taking into account the current effect.
@@
with $C_{1}=\sqrt{a_{1}^{2}+b_{1}^{2}}$, in which $a_{1}$ and $b_{1}$ denote the first sine and cosine Fourier coefficients of the directional spectrum, respectively [57]. Note that alternative and equivalent descriptions for the directional spectrum and spreading exist, such as directional spread $\sigma_{\theta}=\sqrt{2(1+s)}$ arriving from a spectral function $D(\theta)\sim\cos^{2s}{(\theta/2)}$ [58]. The significant wave height $H_{s}$ was computed as the mean of the highest third of waves in each 30-minute sample. To evaluate the effect of current speed $U$ and current direction on major sea state parameters, we selected two sets of samples, for which the angle between peak wave direction and current direction was either $0\pm 10^{\circ}$ (henceforth referred to as ”following current”) or $180\pm 10^{\circ}$ (henceforth referred to as ”opposing current”). Within these ranges, sine values keep below 0.17 and cosine values stay beyond ±$\pm 0.98$, so that the transverse flow is negligible and the axial current is virtually unaffected. References without current, hereafter denoted rest conditions, were defined as samples with $|U/c_{g}|<0.02$, $c_{g}$ being the group velocity of the peak frequency of the wave spectrum. This threshold is consistent with the residual transverse current at current incidence angles of ±$\pm 10^{\circ}$.
In total, the quality-controlled data set contained 4686 samples (2156 with opposing current, 2321 with following current, 209 in rest conditions). Values of $U/c_{g}$ were binned in intervals of $\sim 1/50$, except around zero, where a single bin covers $|U/c_{g}|\leq 0.02$, and for the outermost bins, where data are sparser.
@@
The histograms in figure 4 provide descriptive statistics of all selected 30 minute samples along the entire data set. Figures 4a,d describe the near-symmetrical character of the tidal speeds, whether in dimensional units or normalized by the peak group velocity. Figures 4b,c provide the wave climate of the region, with mild sea characteristics for both significant wave height and zero-crossing period. Most of the data is in intermediate ($k_{p}h=0.3-3.1$, 47% of the dataset) to deep water ($k_{p}h>\pi$, 50% of the dataset, or even 79% corresponding to $k_{p}h\geq 2$) (figure 4e). Adding to the complexity of the sea conditions, the spectrum is significantly directionally spread (figure 4f). Furthermore, the spectrum is very broad (Figure 4h). The sea is composed of waves typically of second order in wave steepness ($0.05\leq\varepsilon\leq 0.09$) [59], albeit smaller fractions of the dataset are also of first and higher orders (figure 4g). To be precise, Stokes expansion of the velocity potential and surface elevation allows us to describe water waves even with fifth-order theory [60] regardless of the numerical value of $\varepsilon$. Nonetheless, in the range $0.05\leq\varepsilon\leq 0.09$ the second-order theory is sufficient. As discussed in the next section, this strong nonlinearity will be key to the physical behavior of the waves. Typically, consideration of the nonlinear regime of the wave steepness and its effect on wave-tide modulation is overlooked in most studies.
III Results
@@
Figure 5 displays the dependence of the main wave parameters on the relative current velocity. Note that there are two ways to normalize the current: either by the current-free group velocity $\bf{c_{g}}$ or the total group velocity $\bf{c_{h}}=\bf{c_{g}}+\bf{U}$ that is affected by the current. Following Longuet-Higgins and Stewart [11], since we are dealing with data whose collected waves are co-linear to the current we can write the normalized speed as either $U/c_{g}=\gamma$ or alternatively $U/c_{h}\equiv\gamma/(1+\gamma)$. One can verify that the modulations in amplitude and wavenumber described by Longuet-Higgins and Stewart [11] can all be expressed as functions of $\gamma$ without further algebra to otherwise have functions of $\gamma/(1+\gamma)$. Thus, we proceed to plot normalized velocities in the direction of motion as $U/c_{g}$. Negative current velocity refers to a current opposing the peak wave direction. Only the mean wave steepness (figure 5a) displays a clear asymmetry between forward and opposing currents. In spite of the wide absolute confidence ranges, the size of the samples is quite large, typically with 100,000 entries per bin. As such, pairwise Welch’s $t$-tests [61] show a highly significant difference ($p<10^{-3}$) between bins corresponding to forward and opposing currents of the same speed, ±$\pm U/c_{g}$, for all pairs of bins.
Other variables display a much more symmetric behavior. The anticipated increase in significant wave height due to an opposing current [1, 10] is observed in our data (figure 5b), but restricted to $|U/c_{g}|\leq 0.05$. Beyond, this trend is inverted and $H_{s}$ decreases again. This decrease might be due to either wave blocking [62] or increasing dissipation due to high mean steepness, although it seems to appear earlier than expected by Toffoli et al. [63]. The behavior is comparable to the effect of following currents: $H_{s}$ increases for mild current speeds, unlike the expected behavior of regular waves [24]. As in the case of an opposing stream, the significant wave height decreases for $U/c_{g}>0.05$, likely due to wave dissipation.
@@
From the classical literature [10], we know that the dispersion relation of narrow-banded, unidirectional, and linear waves is affected by the presence of a current: The dispersion relation $\omega_{0}^{2}=gk_{0}$ becomes $(\omega\pm Uk)^{2}=gk_{0}$. Bearing in mind that in deep water $\omega/k=c_{p}=2c_{g}$, with $c_{p}$ denoting the phase speed, one finds the modulation of the wave number to be of the order of,
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\frac{k}{k_{0}}\approx\frac{1}{[1\pm(U/2c_{g})]^{2}}\approx 1\mp(U/c_{g})+3/4(U/c_{g})^{2}\quad.}$ (4)
When $|U/c_{g}|\ll 0.1$, this modulation is expected to be strongly asymmetrical, however larger values of $U/c_{g}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}=\gamma}$ will diminish such asymmetry, as the quadratic term increases and tends to dominate. A less simplistic computation can be performed for non-homogeneous currents, and Longuet-Higgins and Stewart [11] showed that polynomials with order of $\gamma$ and $\gamma^{2}$ will appear. Furthermore, the wave theory for steeper waves leads to a strengthening of the quadratic term [64, 65, 66], in addition to shear terms [67, 23]. Besides, from the dynamic point of view, Ho et al. [43] showed that the relevant parameter is the phase speed of the current relative to the wave group velocity, so that the interaction between a following current and a comparatively slow wave field in the fixed frame appears to be an interaction between an opposing current and waves in the reference frame of the tide.